OpenStudy (anonymous):

John travelled from Greenviller to Hanover at an average speed of 64 km/hr, By travelling at an average rate of 78 km/hr, he could have arrived 10 minutes earlier. How far is it from Greenville to Hanover?

OpenStudy (anonymous):

let t = time in mins, d = distance b/w two places use time = distance/speed eqn 1: t = d/64 eqn 2: t-10=d/78 solve the linear system for d

OpenStudy (anonymous):

hoow would you do thaat ?

OpenStudy (anonymous):

o.o why not? there's a formula for speed. (Speed = Distance/time). We just rearrange it for time = distance/speed

OpenStudy (anonymous):

sorry hehe I'm in class right now... the first 64km/h shows the speed so we got eqn 1 the 78km/h becomes eqn 2 but the time t must be 10mins less

OpenStudy (anonymous):

so you would divide the distance by the speed ? what would you get as an answer ? im confusedd ..

OpenStudy (anonymous):

Sorry, what grade is this? I used the linear system method to solve this. If you haven't learn it yet, then obviously you couldn't do it this way.